# Transformers as Unsupervised Learning Algorithms: A study on Gaussian Mixtures

### v1

@@ C.3.2 Convergence results for empirical-EM algorithm
Now taking $\delta=\varepsilon$, $A={\left((1-a)\pi_{\min}\right)}^{-1}$ in Lemma C.4 and $\delta=\varepsilon/K$, $A=(R_{\max}+a)^{2}$ in Lemma C.5, we have $\|\hat{u}_{\ell}-u_{\ell}\|_{\infty}\leq 3\varepsilon/2$, hence
@@
Then by the construction in Section C.2, we have
@@
For the term $\|\hat{\mu}_{i}^{+}-\mu_{i}^{+}\|$, we can calculate it similar to the proof of Theorem C.6. First, we recall that with probability at least $21-\delta/$ 2,
@@
Similarly, for $\hat{w}_{\ell,i}$, we can also get(just calculate again) that with probability at least $21-\delta/$ 2,
@@
Then following the same argument in Theorem C.6, taking $A=\exp{\left(\tilde{O}(K_{0}R_{\max}(R_{\max}+d_{0}))\right)}$ and and $\delta=\varepsilon$ in Lemma C.4, we have also
@@
Combining Eq. 21 and Eq. 22, Eq. 20 is proved. The parameter bounds can be directly computed by the construction in Section C.2, Lemma C.4, Lemma C.5 and the parameter $A$, $\delta$ taken in the proof.
C.4 Proof of Theorem C.1
First, by Theorem C.6 and the first condition in Theorem C.1 , there exist a $2$-layer transformer ${\rm TF}_{{\boldsymbol{\Theta}}_{1}}$ such that
@@
holds with probability at least $21-\delta/$ 2. Then using Theorem C.3, Eq. 23 and the second condition in Theorem C.1, there $2$-layer transformer ${\rm TF}_{{\boldsymbol{\Theta}}_{2}}$ such that
@@
uniformly holds with probability at least $21-\delta/$ 2. Denote as ${\boldsymbol{\Theta}}_{2}^{L}=\cup_{\ell\in[L]}{\boldsymbol{\Theta}}_{2}$. Thus, for any $L\in\mathbb{N}$, by reduction, we have
@@
uniformly holds with probability at least $21-\delta/$ 2. Combine all things together, we have, for any $L\in\mathbb{N}$,
@@
holds with probability at least $1-\delta$ (Note that the definitions of $\epsilon{\left(\cdot\right)}$ in Theorem C.7 and Theorem C.1 differ slightly). The parameter bounds can be directly computed by Theorem C.6 and Theorem C.7. The theorem is proved.
Appendix D Formal statement of Theorem 2 and proofs
Following Appendix C, we implement $\operatorname{Readin}$ as an identity transformation and define $\operatorname{Readout}$ to extract targeted matrix elements hence they are both fixed functions.
D.1 Formal statement of Theorem 2
In this section, we give the formal statement of Theorem 2. First, we need to introduce the embeddings of the transformer. Let ${\mathbf{T}}$ be the matrix representation of the cubic tensor $T$, which is
@@ D.1 Formal statement of Theorem 2
where $\overline{{\mathbf{t}}}_{i}\in\mathbb{R}^{d_{0}^{2}}$ is defined as Eq. 24, $\overline{v}^{(0)}={\left[v^{(0)\top},{\mathbf{0}}_{d_{0}-d}^{\top}\right]}^{\top}\in\mathbb{R}^{d_{0}}$, ${\mathbf{e}}_{i}\in\mathbb{R}^{d_{0}}$ denotes the $i$-th standard unit vector in $\mathbb{R}^{d_{0}}$. We choose $D=O(d_{0}^{2})$ and $\tilde{D}=D-d_{0}^{2}-2d_{0}-2$ to get the encoding above. Then we give a rigorous definition of ReLU-activated transformer following [4].
@@
The MLP layer is the same as Definition 2. The ReLU-activated transformer is defined as follows.
@@
Similar to Section 4.1, We consider the following function class of transformer.
@@
Now we can give the formal statement of Theorem 2.
Theorem D.8 (Formal version of Theorem 2).
There exists a transformer ${\rm TF}_{\boldsymbol{\Theta}}$ with ReLU activation such that for any $d\leq d_{0}$, $T\in\mathbb{R}^{d\times d\times d}$ and $v^{(0)}\in\mathbb{R}^{d}$, given the encoding Eq. 25, ${\rm TF}_{\boldsymbol{\Theta}}$ implements $L$ steps of Eq. 4 exactly. Moreover, ${\rm TF}_{\boldsymbol{\Theta}}$ falls within the class ${\mathcal{F}}$ with parameters satisfying:
@@
Remark D.1.
In fact, Theorem D.8 is also hold for attention-only transformers since the MLP layer do not use in the proof. To do that, we only need to add another head in every odd attention layer to clean the terms ${\{d\overline{v}_{i}\}}$. For details, see the proof.
Remark D.2.
@@
Remark D.3.
The use of the ReLU activation function here is primarily for technical convenience and does not alter the fundamental nature of the attention mechanism. Several studies have demonstrated that transformers with ReLU-based attention perform comparably to those using softmax attention[41, 4, 16].
D.2 Proof of Theorem D.8
@@
$\widetilde{{\mathbf{h}}}_{i}^{(1)}={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(1)}}^{R}{\left({\mathbf{H}}^{0}\right)}\right]}_{:,i}={\mathbf{h}}_{i}^{(0)}+\frac{1}{d}\sum_{j=1}^{d}{\left({\left({\mathbf{Q}}^{(1)}{\mathbf{h}}_{i}^{0}\right)}^{\top}{\left({\mathbf{K}}^{(1)}{\mathbf{h}}_{j}^{0}\right)}\right)}{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{0}={\mathbf{h}}_{i}^{(0)}+\begin{bmatrix}{\mathbf{0}}_{d_{0}^{2}}\\ {\mathbf{0}}_{d_{0}}\\ 0\\ 0\\ d\overline{v}^{(0)}_{i}\\ {\mathbf{0}}\end{bmatrix}=\begin{bmatrix}\overline{{\mathbf{t}}}_{i}\\ \overline{v}^{(0)}\\ 1\\ d\\ d\overline{v}^{(0)}_{i}\\ {\mathbf{0}}\end{bmatrix},~{}i\in[d].$
@@
$={\mathbf{h}}_{i}^{(1)}+\sum_{m=1}^{d_{0}}\frac{1}{d}\sum_{j=1}^{d}{\left({\left({\mathbf{Q}}^{(2)}_{m}{\mathbf{h}}_{i}^{(1)}\right)}^{\top}{\left({\mathbf{K}}^{(2)}_{m}{\mathbf{h}}_{j}^{(1)}\right)}\right)}{\mathbf{V}}^{(2)}_{m}{\mathbf{h}}_{j}^{(1)}+\frac{1}{d}\sum_{j=1}^{d}{\left({\left({\mathbf{Q}}^{(2)}_{d_{0}+1}{\mathbf{h}}_{i}^{(1)}\right)}^{\top}{\left({\mathbf{K}}^{(2)}_{d_{0}+1}{\mathbf{h}}_{j}^{(1)}\right)}\right)}{\mathbf{V}}^{(2)}_{d_{0}+1}{\mathbf{h}}_{j}^{(1)}$
@@
Appendix E More on empirical studies
E.1 More on experimental setups
@@
We adopt a Mamba2 [8] model comprising $12$-layers and $128$-dimensional hidden states, with the rest hyper-parameters chosen so as to approximately match the number of a $12$-layer transformer with $128$-dimensional hidden states. As the Mamba series of models are essentially recurrent neural networks (RNNs), we tested two different kinds of $\operatorname{Readout}$ design with either (i). the attentive pooling module as used in the case of transformer backbone and (ii). a more natural choice of using simply the last hidden state to decode all the estimates, as RNNs compress input information in an ordered fashion. We observe from our empirical investigations that using attentive pooling yields better performance even with a Mamba2 backbone. The other training configurations are cloned from those in TGMM experiments with transformer backbones.
@@
Our framework is built upon PyTorch [37] and transformers [46] libraries, which are open-source software released under BSD-style  [^2] https://github.com/pytorch/pytorch/blob/master/LICENSE and Apache license  [^3] https://github.com/huggingface/transformers/blob/main/LICENSE. The code implementations will be open-sourced after the reviewing process of this paper. All the experiments are conducted using $8$ NVIDIA A100 GPUs with $80$ GB memory each.
E.2 A complete report regarding different evaluation metrics
@@
The results are reported in figure 10, 11, 12, 13 and 14, respectively. According to the evaluations, the learned TGMM models show comparable clustering accuracy against the spectral algorithm and outperform EM algorithm when $K>2$ across all comparisons. Regarding the log-likelihood metric, TGMM demonstrates comparable performance with the other two classical algorithms in comparatively lower dimensional cases. i.e., $d\in{\{2,8\}}$, but underperforms both baselines in larger dimensional problems. We conjecture that is might be due to the fact that EM algorithm is essentially a maximum-likelihood algorithm [10], while the TGMM estimation objective (2) is not explicitly related to likelihood-based training.
@@
Figure 11: Assessments of TGMM under test-time task distribution shifts I: A line with $N_{0}^{\text{train}}\rightarrow N^{\text{test}}$ draws the performance of a TGMM model trained over tasks with sample size randomly sampled in $[N_{0}^{\text{train}}/2,N_{0}^{\text{train}}]$ and evaluated over tasks with sample size $N^{\text{test}}$. We can view the configuration $128\rightarrow 128$ as an in-distribution test and rest as out-of-distribution tests.
@@
E.3 On the impact of inference-time sample size $N$
@@
E.4 On the impact of backbone scale
The scaling phenomenon is among the mostly discussed topics in modern AI, as choosing a suitable scale is often critical to the performance of transformer-based architectures like LLMs. In this section we investigate the scaling properties of TGMM via comparing performances produced by varying sizes of backbones that differ either in per-layer width (i.e., the dimension of attention embeddings) or in the total number of layers $L$. With the rest hyper-parameters controlled to be the same as those in section 3.1. The results are reported in three metrics and summarized in figure 16 and figure 17, respectively. According to these investigations, while in general a larger-sized backbone yields slightly better performance as compared to smaller ones. The performance gaps remain mild especially for tasks with relative lower complexity, i.e., $K=2$. Consequently, even a $3$-layer transformer backbone is able to achieve non-trivial learning performance for solving isotropic GMMs, a phenomenon that was also observed in a recent work [17].

### v2

@@ F.3.2 Convergence results for empirical-EM algorithm
Now taking $\delta=\varepsilon$, $A={\left((1-a)\pi_{\min}\right)}^{-1}$ in Lemma˜F.4 and $\delta=\varepsilon/K$, $A=(R_{\max}+a)^{2}$ in Lemma˜F.5, we have $\|\hat{u}_{\ell}-u_{\ell}\|_{\infty}\leq 3\varepsilon/2$, hence
@@
Then by the construction in Section˜F.2, we have
@@
For the term $\|\hat{\mu}_{i}^{+}-\mu_{i}^{+}\|$, we can calculate it similar to the proof of Theorem˜F.6. First, we recall that with probability at least $1-\delta/2$,
@@
Similarly, for $\hat{w}_{\ell,i}$, we can also get(just calculate again) that with probability at least $1-\delta/2$,
@@
Then following the same argument in Theorem˜F.6, taking $A=\exp{\left(\tilde{O}(K_{0}R_{\max}(R_{\max}+d_{0}))\right)}$ and and $\delta=\varepsilon$ in Lemma˜F.4, we have also
@@
Combining Eq.˜20 and Eq.˜21, Eq.˜19 is proved. The parameter bounds can be directly computed by the construction in Section˜F.2, Lemma˜F.4, Lemma˜F.5 and the parameter $A$, $\delta$ taken in the proof.
F.4 Proof of Theorem˜F.1
First, by Theorem˜F.6 and the first condition in Theorem˜F.1 , there exist a $2$-layer transformer ${\rm TF}_{{\boldsymbol{\Theta}}_{1}}$ such that
@@
holds with probability at least $1-\delta/2$. Then using Theorem˜F.3, Eq.˜22 and the second condition in Theorem˜F.1, there $2$-layer transformer ${\rm TF}_{{\boldsymbol{\Theta}}_{2}}$ such that
@@
uniformly holds with probability at least $1-\delta/2$. Denote as ${\boldsymbol{\Theta}}_{2}^{L}=\cup_{\ell\in[L]}{\boldsymbol{\Theta}}_{2}$. Thus, for any $L\in\mathbb{N}$, by reduction, we have
@@
uniformly holds with probability at least $1-\delta/2$. Combine all things together, we have, for any $L\in\mathbb{N}$,
@@
holds with probability at least $1-\delta$ (Note that the definitions of $\epsilon{\left(\cdot\right)}$ in Theorem˜F.7 and Theorem˜F.1 differ slightly). The parameter bounds can be directly computed by Theorem˜F.6 and Theorem˜F.7. The theorem is proved.
Appendix G Formal statement of Theorem˜2 and proofs
Following Appendix˜F, we implement $\operatorname{Readin}$ as an identity transformation and define $\operatorname{Readout}$ to extract targeted matrix elements hence they are both fixed functions.
G.1 Formal statement of Theorem˜2
In this section, we give the formal statement of Theorem˜2. First, we need to introduce the embeddings of the transformer. Let $\mathbf{T}$} be the matrix representation of the cubic tensor $T$, which is
@@ G.1 Formal statement of Theorem˜2
where $\overline{{\mathbf{t}}}_{i}\in\mathbb{R}^{d_{0}^{2}}$ is defined as Eq.˜23, $\overline{v}^{(0)}={\left[v^{(0)\top},{\mathbf{0}}_{d_{0}-d}^{\top}\right]}^{\top}\in\mathbb{R}^{d_{0}}$, ${\mathbf{e}}_{i}\in\mathbb{R}^{d_{0}}$ denotes the $i$-th standard unit vector in $\mathbb{R}^{d_{0}}$. We choose $D=O(d_{0}^{2})$ and $\tilde{D}=D-d_{0}^{2}-2d_{0}-2$ to get the encoding above. Then we give a rigorous definition of ReLU-activated transformer following Bai et al. (2023).
@@
The MLP layer is the same as Definition˜5. The ReLU-activated transformer is defined as follows.
@@
Similar to Section˜4.1, We consider the following function class of transformer.
@@
Now we can give the formal statement of Theorem˜2.
Theorem G.8 (Formal version of Theorem˜2).
There exists a transformer ${\rm TF}_{\boldsymbol{\Theta}}$ with ReLU activation such that for any $d\leq d_{0}$, $T\in\mathbb{R}^{d\times d\times d}$ and $v^{(0)}\in\mathbb{R}^{d}$, given the encoding Eq.˜24, ${\rm TF}_{\boldsymbol{\Theta}}$ implements $L$ steps of Eq.˜3 exactly. Moreover, ${\rm TF}_{\boldsymbol{\Theta}}$ falls within the class $\mathcal{F}$} with parameters satisfying:
@@
Remark G.1.
In fact, Theorem˜G.8 is also hold for attention-only transformers since the MLP layer do not use in the proof. To do that, we only need to add another head in every odd attention layer to clean the terms $\{d\overline{v}_{i}\}$}. For details, see the proof.
Remark G.2.
@@
Remark G.3.
The use of the ReLU activation function here is primarily for technical convenience and does not alter the fundamental nature of the attention mechanism. Several studies have demonstrated that transformers with ReLU-based attention perform comparably to those using softmax attention(Shen et al., 2023; Bai et al., 2023; He et al., 2025a).
G.2 Proof of Theorem˜G.8
@@
$\widetilde{{\mathbf{h}}}_{i}^{(1)}$ $={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(1)}}^{R}{\left({\mathbf{H}}^{0}\right)}\right]}_{:,i}={\mathbf{h}}_{i}^{(0)}+\frac{1}{d}\sum_{j=1}^{d}{\left({\left({\mathbf{Q}}^{(1)}{\mathbf{h}}_{i}^{0}\right)}^{\top}{\left({\mathbf{K}}^{(1)}{\mathbf{h}}_{j}^{0}\right)}\right)}{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{0}$
$={\mathbf{h}}_{i}^{(0)}+\begin{bmatrix}{\mathbf{0}}_{d_{0}^{2}}\\ {\mathbf{0}}_{d_{0}}\\ 0\\ 0\\ d\overline{v}^{(0)}_{i}\\ {\mathbf{0}}\end{bmatrix}=\begin{bmatrix}\overline{{\mathbf{t}}}_{i}\\ \overline{v}^{(0)}\\ 1\\ d\\ d\overline{v}^{(0)}_{i}\\ {\mathbf{0}}\end{bmatrix},~i\in[d].$
@@
$={\mathbf{h}}_{i}^{(1)}+\sum_{m=1}^{d_{0}}\frac{1}{d}\sum_{j=1}^{d}{\left({\left({\mathbf{Q}}^{(2)}_{m}{\mathbf{h}}_{i}^{(1)}\right)}^{\top}{\left({\mathbf{K}}^{(2)}_{m}{\mathbf{h}}_{j}^{(1)}\right)}\right)}{\mathbf{V}}^{(2)}_{m}{\mathbf{h}}_{j}^{(1)}$
$\qquad+\frac{1}{d}\sum_{j=1}^{d}{\left({\left({\mathbf{Q}}^{(2)}_{d_{0}+1}{\mathbf{h}}_{i}^{(1)}\right)}^{\top}{\left({\mathbf{K}}^{(2)}_{d_{0}+1}{\mathbf{h}}_{j}^{(1)}\right)}\right)}{\mathbf{V}}^{(2)}_{d_{0}+1}{\mathbf{h}}_{j}^{(1)}$
@@
Appendix H More on empirical studies
H.1 More on experimental setups
@@
We adopt a Mamba2 Dao and Gu (2024) model comprising $12$-layers and $128$-dimensional hidden states, with the rest hyper-parameters chosen so as to approximately match the number of a $12$-layer transformer with $128$-dimensional hidden states. As the Mamba series of models are essentially recurrent neural networks (RNNs), we tested two different kinds of $\operatorname{Readout}$ design with either (i). the attentive pooling module as used in the case of transformer backbone and (ii). a more natural choice of using simply the last hidden state to decode all the estimates, as RNNs compress input information in an ordered fashion. We observe from our empirical investigations that using attentive pooling yields better performance even with a Mamba2 backbone. The other training configurations are cloned from those in TGMM experiments with transformer backbones.
@@
Our framework is built upon PyTorch Paszke et al. (2019) and transformers Wolf et al. (2020) libraries, which are open-source software released under BSD-style  [^2] https://github.com/pytorch/pytorch/blob/master/LICENSE and Apache license  [^3] https://github.com/huggingface/transformers/blob/main/LICENSE. The code implementations will be open-sourced after the reviewing process of this paper. All the experiments are conducted using $8$ NVIDIA A100 GPUs with $80$ GB memory each.
H.2 A complete report regarding different evaluation metrics
@@
The results are reported in figure 10, 11, 12, 13 and 14, respectively. According to the evaluations, the learned TGMM models show comparable clustering accuracy against the spectral algorithm and outperform EM algorithm when $K>2$ across all comparisons. Regarding the log-likelihood metric, TGMM demonstrates comparable performance with the other two classical algorithms in comparatively lower dimensional cases. i.e., $d\in{\{2,8\}}$, but underperforms both baselines in larger dimensional problems. We conjecture that is might be due to the fact that EM algorithm is essentially a maximum-likelihood algorithm Dempster et al. (1977), while the TGMM estimation objective (2) is not explicitly related to likelihood-based training.
@@
Figure 11: Assessments of TGMM under test-time task distribution shifts I: A line with $N_{0}^{\text{train}}\rightarrow N^{\text{test}}$ draws the performance of a TGMM model trained over tasks with sample size randomly sampled in $N_{0}^{\text{train}}/2,N_{0}^{\text{train}}]$ and evaluated over tasks with sample size $N^{\text{test}}$. We can view the configuration $128\rightarrow 128$ as an in-distribution test and rest as out-of-distribution tests.
@@
H.3 On the impact of inference-time sample size NN
@@
H.4 On the impact of backbone scale
The scaling phenomenon is among the mostly discussed topics in modern AI, as choosing a suitable scale is often critical to the performance of transformer-based architectures like LLMs. In this section we investigate the scaling properties of TGMM via comparing performances produced by varying sizes of backbones that differ either in per-layer width (i.e., the dimension of attention embeddings) or in the total number of layers $L$. With the rest hyper-parameters controlled to be the same as those in section 3.1. The results are reported in three metrics and summarized in figure 16 and figure 17, respectively. According to these investigations, while in general a larger-sized backbone yields slightly better performance as compared to smaller ones. The performance gaps remain mild especially for tasks with relative lower complexity, i.e., $K=2$. Consequently, even a $3$-layer transformer backbone is able to achieve non-trivial learning performance for solving isotropic GMMs, a phenomenon that was also observed in a recent work He et al. (2025b).